# Spline Quantile Regression

### v1

@@ Abstract
Quantile regression is a powerful tool capable of offering a richer view of the data as compared to linear-squares regression. Quantile regression is typically performed individually on a few quantiles or a grid of quantiles without considering the similarity of the underlying regression coefficients at nearby quantiles. When needed, an ad hoc post-processing procedure such as kernel smoothing is employed to smooth the estimated coefficients across quantiles and thereby improve the performance of these estimates. This paper introduces a new method, called spline quantile regression (SQR), that unifies quantile regression with quantile smoothing and jointly estimates the regression coefficients across quantiles as smoothing splines. We discuss the computation of the SQR solution as a linear program (LP) using an interior-point algorithm. We also experiment with some gradient algorithms that require less memory than the LP algorithm. The performance of the SQR method and these algorithms is evaluated using simulated and real-world data.
@@ 1 Introduction
We offer an alternative solution, called spline quantile regression (SQR). The SQR method extends the original QR problem at an individual quantile into a function estimation problem across all quatiles. In this problem, the regression coefficients are treated as smooth functions of the quantile level in a space of splines, and the resulting model is fitted to the data jointly on a grid of quantiles. A penalty term is added to the QR cost function to regularize the smoothness of the functional regression coefficients. The SQR method unifies quantile regression with quantile smoothing and provides an estimator of regression coefficients as smoothing splines.
@@ 2 Spline Quantile Regression
Let $\{y_{t}:t=1,\dots,n\}$ be a sequence of $n$ observations of a dependent variable and $\{\mathbf{x}_{t}:t=1,\dots,n\}$ be the corresponding values of a $p$-dimensional regressor. Given an increasing sequence of quantile levels $\{\tau_{\ell}:\ell=1,\dots,L\}\subset(0,1)$, the spline quantile regression (SQR) problem can be stated as
@@
where $\mathcal{F}$ is a functional space spanned by spline basis functions, $\rho_{\tau}(y):=y(\tau-{\cal I}(y<0))$ is the objective function of quantile regression at quantile level $\tau$ (Koenker 2005, p. 5), $c\geq 0$ is a smoothing or penalty parameter, and $\{w_{\ell}\}$ is a sequence of nonnegative constants.
The $L_{1}$-norm of second derivatives is employed in (1) as the roughness measure of $\boldsymbol{\beta}(\cdot)$ in order to retain the LP characteristics of the original QR problem (Koenker 2005). The squared $L_{2}$-norm, which is commonly used for spline smoothing under the least-squares framework (Wahba 1975), is a possible alternative. This $L_{2}$-norm penalty would result in a quadratic program. We choose not to deal with the quadratic program in this paper. The weight sequence $\{w_{\ell}\}$ is introduced in (1) to allow different contributions from different quantiles to the penalty term. For example, taking $w_{\ell}:=0.25/(\tau_{\ell}(1-\tau_{\ell}))$ would limit the influence from high and low tails for their excessive statistical variability.
@@
Note that the SQR problem is different from the problem of quantile smoothing splines considered by Koenker et al. (1994). The latter uses splines to represent nonparametric functions of independent variables. The SQR problem is also different from the problem considered by Andriyana et al. (2014) in which splines are used to represent regression coefficients as functions of time rather than the quantile level.
@@ 3 Linear Program Reformulation
The primal-dual pair given by (11) and (12) are in the canonical form required by the interior-point algorithm of Portnoy and Koenker (1997). This algorithm solves the primal-dual pair jointly by using Newton’s method in which positivity constraints are enforced by log barriers (Koenker and Ng 2005). An implementation of this algorithm as a FORTRAN code rqfnb.f is available in the ‘quanreg’ package. The rq.fit.fnb function invokes this code to compute the solution to the ordinary QR problem. For the SQR problem, it suffices to modify the rq.fit.fnb function using the properly modified input variables.
@@
Because the dual problem for SQR is given by (11), it suffices to replace the response vector $\mathbf{y}$ by $\mathbf{b}$, the design matrix $\mathbf{X}$ by $\mathbf{D}$, and the right-hand-side vector $(1-\tau)\mathbf{X}^{T}{\bf 1}_{n}$ in the equality constraints by $\mathbf{a}$. In addition, we set the initial value of $\boldsymbol{\zeta}$ to $[(1-\tau_{1}){\bf 1}_{n}^{T},\dots,(1-\tau_{L}){\bf 1}_{n}^{T},0.5{\bf 1}_{pL}^{T}]^{T}$. We call the resulting R function rq.fit.fnbs. For solving the SQR problem, the interior-point algorithm turns out to be much more efficient than standard LP solvers, such as the lp function, in terms of computer memory and time.
To develop a data-driven method for selecting the smoothing parameter $c$, we adopt a technique used by Koenker et al.  (1994) (see also Koenker 2005, p. 234) for nonparametric quantile regression at a fixed quantile level. In this technique, the mean objective function of quantile regression at the fitted values is treated as the fidelity measure, similar to the residual sum of squares divided by $n$ in linear regression, and the number of points interpolated (or closely approximated) by the fitted values is treated as the complexity measure, similar to the number of parameters in linear regression. We extend these measures to the SQR problem by using the average values across the quantile levels. The resulting SIC criterion takes the form
${\rm SIC}(c):=n\log\bigg{(}L^{-1}\sum_{\ell=1}^{L}v_{c}(\tau_{\ell})\bigg{)}+(1/2)\log n\,\bigg{(}L^{-1}\sum_{\ell=1}^{L}m_{c}(\tau_{\ell})\bigg{)}$ (14)
where $v_{c}(\tau_{\ell}):=n^{-1}\sum_{t=1}^{n}\rho_{\tau_{\ell}}(y_{t}-\mathbf{x}_{t}^{T}\hat{\boldsymbol{\beta}}(\tau_{\ell}))$ is the fidelity measure at $\tau_{\ell}$, and $m_{c}(\tau_{\ell})$, denoting the number of points closely approximated by the fitted values $\mathbf{x}_{t}^{T}\hat{\boldsymbol{\beta}}(\tau_{\ell})$, is the corresponding complexity measure. By the same token, we obtain two variants of the SIC criterion: i.e., the AIC criterion where $(1/2)\log n$ in (14) is replaced by 2, and the BIC criterion where $(1/2)\log n$ is replaced by $\log n$. Clearly, BIC always imposes a heavier penalty on the complexity than SIC, and, for sufficiently large $n$ ($n\geq 55$ so that $(1/2)\log n>2$), SIC imposes a heavier penalty than AIC. As a result, BIC produces the smoothest estimates and AIC the least smooth estimates.
@@ 4 Examples
To demonstrate the performance of the SQR method, we present some examples in this section with real and simulated data.
@@
Our first example is the Engel food expenditure data that comes with the ‘quantreg’ package (Koenker 2005, pp. 300–302). In this example, the household expenditure on food, $y$, is assumed to obey a quantile regression model
$Q_{y|x}(\tau)=\beta_{1}(\tau)+\beta_{2}(\tau)\,(x-\mu),$
@@
where $x$ is the household income with mean $\mu$. The data set contains $n=235$ records.
@@
By following Koenker (2005), we estimate the coefficient functions $\beta_{1}(\cdot)$ and $\beta_{2}(\cdot)$ by QR and SQR on the quantile grid $\{0.08,0.09,\dots,0.98\}$ with $\mu$ replaced by the sample mean of income. Figure 1 shows these estimates as functions of $\tau$. The confidence band is constructed according to the asymptotic normality of QR estimates under the non-iid settings (Koenker 2005, p. 34), as implemented by the summary.rq function in the R package ‘quantreg’ with the option se=‘nid’. In comparison with the QR estimates (open circles in Figure 1), the SQR estimates appear less noisy across quantiles. The SQR estimates from different criteria for smoothing parameter selection do not differ dramatically. But a closer inspection reveals that the results from BIC contain fewer zigzags than those from AIC, especially for estimating $\beta_{2}(\cdot)$.
@@
Figure 1: SQR estimates of the functional coefficients $\beta_{1}(\cdot)$ (left) and $\beta_{2}(\cdot)$ (right) for the Engel food expenditure data. Smoothing parameters are selected by AIC (top), SIC (middle), and BIC (bottom). Dashed lines depict a 95% pointwise confidence band. Open circles depict the QR estimates.
@@
Our second example is the time series of yearly sunspot numbers (Figure 4). This series was examined in Li (2012; 2014) for its periodicity through the quantile periodogram shown in the left panel of Figure 4. The quantile periodogram is an alternative to the conventional periodogram but derived from trigonometric quantile regression. It is defined as $(n/4)[\hat{\beta}_{2}^{2}(\tau,\omega)+\hat{\beta}_{3}^{2}(\tau,\omega)]$, where $\hat{\beta}_{2}(\tau,\omega)$ and $\hat{\beta}_{3}(\tau,\omega)$ are given by QR using the regressor $\mathbf{x}_{t}(\omega):=[1,\cos(\omega t),\sin(\omega t)]^{T}$, where h $\omega$ takes values in the set of Fourier frequencies in interval $(0,\pi)$. As a bivariate function of $\omega$ and $\tau$, the quantile periodogram in the left panel of Figure 4 suggests that the sunspot nubmers not only have the well-known cycle of 11 years, but the magnitude of this cycle decreases with the quantile level, i.e., stronger at higher quantiles and weaker at lower quantiles.
@@
Figure 3: Quantile periodogram of yearly sunspot numbers constructed by trigonometric quantile regression using QR (left) and SQR with AIC (right).
@@
Figure 4: Quantile periodogram, constructed by SQR with AIC, of yearly sunspot numbers at 11 year and 103 year frequencies. Open circles depict the result from QR.
@@
Instead of QR, the trigonometric quantile regression can be performed by SQR. The resulting quantile periodogram is shown in the right panel of Figure 4. To construct this quantile periodogram, a single smoothing parameter is employed in the trigonometric SQR problems for all frequencies. It is chosen by minimizing the average AIC across the frequencies. Needless to say, one may also choose to employ different smoothing parameters for different frequencies to achieve greater flexibility but with higher computational burden and statistical variability.
@@
Compared to the QR-based quantile periodogram, the SQR-based quantile periodogram appears smoother across quantiles, not only near the peaks but also in the background. The difference between these periodograms can be further appreciated by inspecting the cross-section plot in Figure 4, where the quantiles periodograms are shown as functions of the quantile level at two peak frequencies corresponding to the 11 year and 103 year cycles. The effect of smoothing is most notable in the lower curve for quantile levels higher than 0.75.
Our third example is a simulated data set of time series based on a so-called quantile autoregressive (QAR) process which satisfies
$y_{t}=a_{0}(U_{t})+a_{1}(U_{t})\,y_{t-1}.$ (15)
In this model, $\{U_{t}\}$ is a sequence of i.i.d. $U(0,1)$ random variables, $a_{0}(u)$ takes the form of the quantile function of $N(0,0.4^{2})$, $a_{1}(u)$ is a piecewise-linear function $0.85+0.1u+0.25(u-0.5){\cal I}(u>0.5)$. This QAR model is a slight modification of the example in Koenker (2005, p. 262) where $a_{0}(u)$ is the quantile function of $N(0,1)$ and $a_{1}(u)$ is a linear function $0.85+0.25u$.
Figure 6: SQR estimates of $a_{0}(\cdot)$ (left) and $a_{1}(\cdot)$ (right) in (15) from the time series shown in Figure 6, where the smoothing parameter is chosen by BIC. Open circles depict the QR estimates. Dashed lines depict the true values. Absolute errors are 0.027 and 0.015 for SQR, The corresponding errors for QR are 0.026 and 0.024.
Figure 6 shows a series simulated from this QAR model ($n=200)$. Figure 6 shows the SQR estimates of the functional coefficients obtained from this series on the grid of quantile levels $\{0.05,0.06,\dots,0.96\}$, with the smoothing parameter, spar = 1.062, selected by BIC. To measure the accuracy of these estimates, we employ the absolute error which is defined as the average of absolute residuals across the quantile levels. For the SQR estimates in Figure 6, the absolute errors are 0.027 and 0.015, respectively, The corresponding errors of the QR estimates are 0.026 and 0.024. When AIC is used instead of BIC for SQR (not shown), the smoothing parameter reduces to spar = 0.835, and the absolute errors become 0.025 and 0.020. The corresponding results from SIC (not shown) yield absolute errors 0.026 and 0.019 with spar = 0.864. Clearly, a larger smoothing parameter benefits the estimate for the piecewise-linear function $a_{1}(\cdot)$, which is consistent with the underlying assumption of SQR, but may degrade the result for the nonlinear function $a_{0}(\cdot)$, which can only be approximated by piecewise-linear functions.
Figure 7: Mean absolute error (MAE) for estimating the functional coefficient $a_{0}(\cdot)$ (left) and $a_{1}(\cdot)$ (right) in the QAR model (15) by SQR with different values of the smoothing parameter spar. NA stands for the ordinary QR. Results are based on 1000 Monte Carlo runs with $n=200$.
This finding is confirmed by a more comprehensive analysis of the estimation accuracy based on 1000 Monte Carlo runs. The result of this analysis is shown in Figure 7, where the mean absolute error (MAE) is plotted against the smoothing parameter spar for estimating $a_{0}(\cdot)$ and $a_{1}(\cdot)$, respectively. As can be seen, the SQR estimates outperform the QR estimates for a range of smoothing parameter values, especially for estimating $a_{1}(\cdot)$. The MAE increases when the smoothing parameter is too large, especially for estimating $a_{0}(\cdot)$. The best spar for estimating $a_{0}(\cdot)$ is around 0.7 and the best spar for estimating $a_{1}(\cdot)$ is around 1.1. The total MAE for estimating both functions (not shown) is minimized when spar equals 0.9. The results in Table 1 further demonstrate that the data-driven criteria work as expected in selecting a suitable smoothing parameter. The heavier penalty in BIC benefits the estimation of $a_{1}(\cdot)$, whereas the lighter penalty in AIC benefits the estimation of $a_{0}(\cdot)$.
Table 1: Mean Absolute Error of QR and SQR Estimates for the QAR Model (15)
Estimation of $a_{0}(\cdot)$ Estimation of $a_{1}(\cdot)$
$n$ QR SQR-AIC SQR-SIC SQR-BIC QR SQR-AIC SQR-SIC SQR-BIC
200 0.0377 0.0367 0.0369 0.0378 0.0374 0.0353 0.0351 0.0349
500 0.0221 0.0219 0.0221 0.0226 0.0208 0.0195 0.0194 0.0193
Results are based on 1000 Monte Carlo run.
@@
Our final example is the US infant birth weight data from the National Bureau of Economic Research (NBER)333 https://www.nber.org/research/data/vital-statistics-natality-birth-data for year 2022. This example follows the method of Koenker (2005, p. 20) by extracting records for black or white mothers between ages 18 and 45. The study in Koenker (2005) is based on a data set for year 1997 and included 16 explanatory variables. We simplify the study in this example by considering only 8 explanatory variables for the birth weight (in grams): infant’s sex (0 for Girl, 1 for Boy), mother’s race (0 for White, 1 for Black), age (18–45), and marital status (0 for Unmarried, 1 for Married), and prenatal medical care named Precare (0 if the first visit in the first trimester of the pregnancy, 1 if no prenatal visit, 2 if the first visit in the second trimester, 3 if the first visit in the last trimester). For simplicity, mother’s education level, smoking status, and weight gain during pregnancy are not used in the model. As in Koenker (2005), we use a quadratic function to represent the effect of age.
@@
Figure 8 depict the estimated functional coefficients by QR and SQR on the quantile grid $\{0.05,0.06,\dots,0.95\}$. These estimates are obtained from a random sample of $n=50000$ records. The 90% pointwise confidence bands (shaded grey areas) for both types of estimates are constructed using the estimated standard errors of the QR estimates under the iid assumption (se=‘iid’ in summary.rq). As in Koenker (2005, p. 21), the dashed and dotted horizontal lines depict the ordinary least-squares estimate of the mean effect and a 90% confidence interval of the estimate. The QR estimates in Figure 8(a) are very similar to those in Koenker (2005, p. 21). The SQR estimates in Figure 8(b) present the quantile-dependent effect of the explanary variables with better-defined and less-noisy patterns.

### v2

@@ Abstract
Quantile regression is a powerful tool capable of offering a richer view of the data as compared to least-squares regression. Quantile regression is typically performed individually on a few quantiles or a grid of quantiles without considering the similarity of the underlying regression coefficients at nearby quantiles. When needed, an ad hoc post-processing procedure such as kernel smoothing is employed to smooth the individually estimated coefficients across quantiles and thereby improve the performance of these estimates. This paper introduces a new method, called spline quantile regression (SQR), that unifies quantile regression with quantile smoothing and jointly estimates the regression coefficients across quantiles as smoothing splines. We discuss the computation of the SQR solution as a linear program (LP) using an interior-point algorithm. We also experiment with some gradient algorithms that require less memory than the LP algorithm. The performance of the SQR method and these algorithms is evaluated using simulated and real-world data.
@@ 1 Introduction
We offer an alternative solution, called spline quantile regression (SQR). The SQR method extends the original QR problem at an individual quantile into a function estimation problem across all quatiles. In this problem, the regression coefficients are treated as smooth functions of the quantile level in a space of splines, and the resulting model is fitted to the data jointly on a grid of quantiles. A penalty term is added to the QR cost function to regularize the smoothness of the functional regression coefficients. The SQR method unifies quantile regression with quantile smoothing and provides an estimator of regression coefficients as smoothing splines. In this paper, we focus on the computation of the SQR solution. We evaluate and demonstrate the SQR solution with simulated and real data.
@@ 2 Spline Quantile Regression
Let $\{y_{t}:t=1,\dots,n\}$ be a sequence of $n$ observations of a dependent variable and $\{\mathbf{x}_{t}:t=1,\dots,n\}$ be the corresponding values of a $p$-dimensional regressor. Under the quantile regression framework (Koenker 2005), it is assumed that the conditional quantile of $y_{t}$ at a quantile level $\tau\in(0,1)$ given $\mathbf{x}_{t}$ can be expressed as
$Q_{y_{t}}(\tau\mid\mathbf{x}_{t})=\mathbf{x}_{t}^{T}\boldsymbol{\beta}(\tau).$ (1)
Given an increasing sequence of quantile levels $\{\tau_{\ell}:\ell=1,\dots,L\}\subset(0,1)$, the standard QR method produces an estimate for each $\boldsymbol{\beta}_{\ell}:=\boldsymbol{\beta}(\tau_{\ell})$ independently by solving
$\hat{\boldsymbol{\beta}}_{\ell}:=\operatorname*{argmin}_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\sum_{t=1}^{n}\rho_{\tau_{\ell}}(y_{t}-\mathbf{x}_{t}^{T}\boldsymbol{\beta})\quad(\ell=1,\dots,L),$ (2)
where $\rho_{\tau}(y):=y(\tau-{\cal I}(y<0))$ denotes the objective function of quantile regression at quantile level $\tau$ (Koenker 2005, p. 5). In typical situations, the quantile levels in $\{\tau_{\ell}\}$ are specified by the user based on the need of the application. In rare occasions, one may want to consider the quantiles that correspond to all distinct quantile regression solutions (Portnoy 1991; Koenker 2005, p. 303) similar to the quantiles associated with the order statistics.
In this paper, we are interested in estimating $\boldsymbol{\beta}(\tau)$ in (2) as a function of $\tau$ in a closed subinterval of $(0,1)$, e.g., $[\epsilon,1-\epsilon]\subset(0,1)$ for some $\epsilon\in(0,0.5)$. We propose to obtain an estimate by solving the spline quantile regression (SQR) problem
@@
where $\mathcal{F}$ is a functional space spanned by spline basis functions, $c\geq 0$ is a smoothing or penalty parameter, and $\{w_{\ell}\}$ is a user-specified sequence of nonnegative constants that allow different contributions from different quantiles to the penalty term.
We employ the $L_{1}$-norm of second derivatives in (3) as the roughness measure of the functional coefficient $\boldsymbol{\beta}(\cdot)$ in order to retain the LP characteristics of the original QR problem (Koenker 2005). This is the same reason offered in Koenker et al. (1994) to justify the $L_{1}$-norm penalty for the spline-based nonparametric regressor in a nonparamtric quantile regression problem. It leads to a piecewise linear function that can be used to approximate any continuous functions when $\max\{|\tau_{\ell+1}-\tau_{\ell}|\}$ is sufficiently small. An altervative to the $L_{1}$-norm is the squared $L_{2}$-norm, which is commonly used for spline smoothing under the least-squares framework (Wahba 1975). This option would result in a quadratic program (QP) instead of an LP. We choose to deal with this problem elsewhere.
@@
The SQR problem in (3) is different from the problem of quantile smoothing splines considered by Koenker et al. (1994). The latter uses splines to represent nonparametric functions of independent variables in form of $Q_{y}(\tau\mid x)=s(x)$ for fixed $\tau$. The SQR problem in (3) is also different from the problems considered by Andriyana et al. (2014) and Kim (2007) in which splines are used to represent regression coefficients as functions of an auxiliary variable rather than functions of the quantile level $\tau$.
@@ 3 Linear Program Reformulation
The primal-dual pair given by (13) and (14) are in the canonical form required by the interior-point algorithm of Portnoy and Koenker (1997). This algorithm solves the primal-dual pair jointly by using Newton’s method in which positivity constraints are enforced by log barriers (Koenker and Ng 2005). An implementation of this algorithm as a FORTRAN code rqfnb.f is available in the ‘quanreg’ package. The rq.fit.fnb function in this package invokes the FORTRAN code to compute the solution to the ordinary QR problem. For the SQR problem, it suffices to modify the rq.fit.fnb function using the properly modified input variables.
@@
Because the dual problem for SQR is given by (13), it suffices to replace the response vector $\mathbf{y}$ by $\mathbf{b}$, the design matrix $\mathbf{X}$ by $\mathbf{D}$, and the right-hand-side vector $(1-\tau)\mathbf{X}^{T}{\bf 1}_{n}$ in the equality constraints by $\mathbf{a}$. In addition, we set the initial value of $\boldsymbol{\zeta}$ to $[(1-\tau_{1}){\bf 1}_{n}^{T},\dots,(1-\tau_{L}){\bf 1}_{n}^{T},0.5{\bf 1}_{pL}^{T}]^{T}$. We call the resulting R function rq.fit.fnb2. For solving the SQR problem, the interior-point algorithm turns out to be much more efficient than standard LP solvers, such as the lp function, in terms of computer memory and time.
To develop a data-driven method for selecting the smoothing parameter $c$, we adopt a technique used by Koenker et al.  (1994) (see also Koenker 2005, p. 234) for nonparametric quantile regression at a fixed quantile level. In this technique, the mean objective function of quantile regression at the fitted values is treated as the fidelity measure, analogous to the standard error of the residuals in linear regression, and the number of points interpolated (or closely approximated) by the fitted values is treated as the complexity measure, similar to the number of parameters in linear regression. We extend the fidelity and complexity measures to the SQR problem. Let $v_{c}(\tau_{\ell}):=n^{-1}\sum_{t=1}^{n}\rho_{\tau_{\ell}}(y_{t}-\mathbf{x}_{t}^{T}\hat{\boldsymbol{\beta}}(\tau_{\ell}))$ be the fidelity measure at $\tau_{\ell}$, and $m_{c}(\tau_{\ell})$, denoting the number of points closely approximated by the fitted values $\mathbf{x}_{t}^{T}\hat{\boldsymbol{\beta}}(\tau_{\ell})$, be the complexity measure at $\tau_{\ell}$. Then, the resulting Bayesian information criterion (BIC), also known as Schwarz information criterion (SIC), takes the form
${\rm BIC}(c):=2n\log\bigg{(}L^{-1}\sum_{\ell=1}^{L}v_{c}(\tau_{\ell})\bigg{)}+\log n\,\bigg{(}L^{-1}\sum_{\ell=1}^{L}m_{c}(\tau_{\ell})\bigg{)}.$ (16)
By the same token, we obtain the AIC criterion which replaces $\log n$ in (16) with 2. As usual, BIC always imposes a heavier penalty on the complexity than AIC for sufficiently large $n$ such that $\log n>2$. As a result, BIC tends to produce a smoother estimate than AIC.
@@ 4 Examples
First, we evaluate the SQR method by a simulated data set of time series based on a so-called quantile autoregressive (QAR) process which satisfies
$y_{t}=a_{0}(U_{t})+a_{1}(U_{t})\,y_{t-1}.$ (17)
In this model, $\{U_{t}\}$ is a sequence of i.i.d. $U(0,1)$ random variables, $a_{0}(u)$ takes the form of the quantile function of $N(0,0.4^{2})$, $a_{1}(u)$ is a piecewise-linear function $0.85+0.1u+0.25(u-0.5){\cal I}(u>0.5)$. This QAR model is a slight modification of the example in Koenker (2005, p. 262) where $a_{0}(u)$ is the quantile function of $N(0,1)$ and $a_{1}(u)$ is a linear function $0.85+0.25u$. The conditional quantile function of $y_{t}$ given $y_{t-1}$ can be written as
$Q_{y_{t}}(\tau\mid y_{t-1})=a_{0}(\tau)+a_{1}(\tau)\,y_{t-1},\quad 0<\tau<1.$
Figure 1 shows a series simulated from this QAR model ($n=200)$.
(a)                                                 (b)
(c)                                                 (d)
Figure 2: SQR estimates of $a_{0}(\cdot)$ and $a_{1}(\cdot)$ in the QAR model (17) from the time series shown in Figure 1. The smoothing parameter is chosen by (a)-(b) AIC and (c)-(d) BIC. Open circles depict the QR estimates. Dashed lines depict the true values.
Figure 2 depicts the SQR estimates and the QR estimates of $a_{0}(\cdot)$ and $a_{1}(\cdot)$ from this series on the quantile grid $\{0.05,0.06,\dots,0.95\}$. For SQR, the smoothing parameter is selected by AIC and BIC. We employ the absolute errors, $L^{-1}\sum_{\ell=1}^{L}|\hat{a}_{j}(\tau_{\ell})-a_{j}(\tau_{\ell})|$ $(j=0,1)$, to measure the accuracy of these estimates. For the SQR estimates with AIC, the errors are both 0.009; for the SQR estimates with BIC, the errors are 0.009 and 0.008, respectively. The corresponding errors of the QR estimates are 0.015 and 0.013. Visual inspection of Figures 2(b) and (c) shows that the SQR with BIC follows the shape of $a_{1}(\cdot)$ more faithfully than the SQR with AIC.
Figure 3: Total mean absolute error for estimating the functional coefficients in the QAR model (17) by SQR with different values of smoothing parameter spar. NA stands for the ordinary QR. Results are based on 1000 Monte Carlo runs with $n=200$.
A more comprehensive analysis of the estimation accuracy is presented in Figure 3. Based on 1000 Monte Carlo runs, Figure 3 shows the total mean absolute error $\text{MAE}_{0}+\text{MAE}_{1}$ against the value of smoothing parameter spar, where
$\text{MAE}_{j}:=E\bigg{\{}L^{-1}\sum_{\ell=1}^{L}(|\hat{a}_{j}(\tau_{\ell})-a_{j}(\tau_{\ell})|\bigg{\}}\quad(j=0,1).$
As can be seen, the SQR estimates outperform the QR estimates for a range of smoothing parameter values, with the best choice being spar = 0.9. Table 1 further demonstrates that the data-driven criteria AIC and BIC are both able to select a smoothing parameter for SQR to outperform QR, with BIC yielding a smaller total MAE for both sample sizes. The improved accuracy of SQR over QR comes largely from estimating $a_{1}(\cdot)$.
Table 1: Mean Absolute Error of SQR Estimates with AIC and BIC for the QAR Model (17)
$\text{MAE}_{0}$ $\text{MAE}_{1}$ Total MAE
$n$ QR SQR-AIC SQR-BIC QR SQR-AIC SQR-BIC QR SQR-AIC SQR-BIC
200 0.0377 0.0367 0.0366 0.0374 0.0356 0.0351 0.0751 0.0723 0.0717
500 0.0221 0.0217 0.0221 0.0208 0.0198 0.0194 0.0429 0.0415 0.0411
Results are based on 1000 Monte Carlo run.
@@
Having validated the performance of the SQR estimator with simulated data, we then present some real-data examples in the remainder of this section.
@@
The first real-data example is the Engel food expenditure data that comes with the ‘quantreg’ package (Koenker 2005, pp. 300–302). It is assume that the household expenditure on food, $y$, obeys a quantile regression model
$Q_{y}(\tau\mid x)=\beta_{1}(\tau)+\beta_{2}(\tau)\,(x-\mu),$
@@
where $x$ is the household income (×$\times 1000$) with mean $\mu$. The data set contains $n=235$ records.
@@
By following Koenker (2005), we estimate the coefficients $\beta_{1}(\cdot)$ and $\beta_{2}(\cdot)$ by QR on the quantile grid $\{0.02,0.03,\dots,0.98\}$ with $\mu$ replaced by the sample mean of income. These estimates are compared with the SQR estimates obtained with automatically selected smoothing parameter by AIC and BIC. The results are shown in Figure 4 as functions of $\tau$. The confidence band for the SQR estimates is constructed according to the asymptotic normality of quantile regression estimates under the non-iid settings (Koenker 2005, p. 34), as implemented by the summary.rq function in the R package ‘quantreg’ with the option se=‘nid’. In comparison with the QR estimates (open circles in Figure 4), the SQR estimates are less noisy and more visually appealing. As expected, BIC produces a “smoother” estimate than AIC due to a heavier penalty on the model complexity. It seems particularly appropriate for BIC to smooth out the wiggly pattern of the QR estimates between $\tau=0.2$ and 0.6, which AIC fails to do.
@@
Figure 4: SQR estimates of the functional coefficients $\beta_{1}(\cdot)$ (left) and $\beta_{2}(\cdot)$ (right) for the Engel food expenditure data. Smoothing parameters are selected by AIC (top) and BIC (bottom). Dashed lines depict a 95% pointwise confidence band. Open circles depict the QR estimates.
@@
The second real-data example is the time series of yearly sunspot numbers (Figure 7). This series was examined in Li (2012; 2014) for its periodicity through the so-called quantile periodogram shown in the left panel of Figure 7. This quantile periodogram is an alternative to the conventional periodogram for spectral analysis of time sereis. It is derived from trigonometric quantile regression instead of Fourier transform. Formally, with $f$ taking values in the set of normalized Fourier frequencies in interval $(0,0.5)$ and $\tau$ taking values in the interval $(0,1)$, the quantile periodogram is defined as $(n/4)[\hat{\beta}_{2}^{2}(f,\tau)+\hat{\beta}_{3}^{2}(f,\tau)]$, where $\hat{\beta}_{2}(f,\tau)$ and $\hat{\beta}_{3}(f,\tau)$ are given by QR using the regressor $\mathbf{x}_{t}(f):=[1,\cos(2\pi ft),\sin(2\pi ft)]^{T}$, As a bivariate function of $f$ and $\tau$, the quantile periodogram in the left panel of Figure 7 suggests that the sunspot nubmers not only have the well-known cycle of 11 years ($f=28/308\approx 0.091$), but the magnitude of this cycle exhibits a generally increasing trend with the increase of quantile level, i.e., the cycle tends to be stronger at higher quantiles and weaker at lower quantiles.
@@
Figure 6: Quantile periodogram of yearly sunspot numbers constructed by trigonometric quantile regression using QR (left) and SQR with BIC (right).
@@
Figure 7: Cross-section of quantile periodogram, constructed by SQR with BIC, of yearly sunspot numbers at 11 year and 103 year frequencies. Open circles depict the result from QR.
@@
Instead of QR, the trigonometric quantile regression can be performed by SQR. The resulting quantile periodogram is shown in the right panel of Figure 7. To construct this quantile periodogram, a single smoothing parameter is employed in the trigonometric SQR problems for all frequencies. It is chosen by minimizing the average BIC across the frequencies. Needless to say, one may also choose to employ different smoothing parameters for different frequencies to achieve greater flexibility but with higher computational burden and statistical variability.
@@
Compared to the QR-based quantile periodogram, the SQR-based quantile periodogram appears smoother across quantiles, not only around the peak frequencies but also in the background. The difference between these periodograms can be further appreciated by inspecting the cross-section plot in Figure 7, where the quantile periodograms are shown as functions of $\tau$ at two peak frequencies corresponding to the 11 year cycle $(f=28/308)$ and the 103 year cycle $(f=2/308)$. Compared to the QR estimtes, the SQR estimates depict a better-defined less-noisy trend which grows steadily with $\tau$. The effect of smoothing is most notable at higher quantiles, especially in the curve of the 103 year cycle.
@@
The final real-data example is the US infant birth weight data from the National Bureau of Economic Research (NBER)333 https://www.nber.org/research/data/vital-statistics-natality-birth-data for year 2022. A similar data set from year 1997 was used in Koenker (2005, p. 20) to demonstrate the usefulness of the quantile regression coefficients as functions of $\tau$. By following this study, we extract the data records for black or white mothers between ages 18 and 45. The study in Koenker (2005, p. 20) included 16 explanatory variables for the birth weight (in grams). We simplify the study by considering only 8 explanatory variables: infant’s sex (0 for Girl, 1 for Boy), mother’s race (0 for White, 1 for Black), age (18–45), and marital status (0 for Unmarried, 1 for Married), and prenatal medical care named Precare (0 if the first visit in the first trimester of the pregnancy, 1 if no prenatal visit, 2 if the first visit in the second trimester, 3 if the first visit in the last trimester). For simplicity, mother’s education level, smoking status, and weight gain during pregnancy are not used in the model. As in Koenker (2005, p. 22), we use a quadratic function to represent the effect of age.
@@
By following Koenker (2005, p. 21), we compute the QR estimates on the quantile grid $\{0.05,0.06,\dots,0.95\}$. These estimates are obtained from a random sample of $n=50000$ records. The corresponding SQR estimates are computed with spar $=-1$. Figure 8 shows the QR and SQR estimates as functions of $\tau$, The 90% pointwise confidence bands (shaded grey areas) for both types of estimates are constructed using the estimated standard errors of quantile regression estimates under the iid assumption (se=‘iid’ in summary.rq). As in Koenker (2005, p. 21), the dashed and dotted horizontal lines depict the ordinary least-squares estimate of the mean effect and a 90% confidence interval of the estimate. The QR estimates in Figure 8(a) are very similar to those in Koenker (2005, p. 21). The SQR estimates in Figure 8(b) present the quantile-dependent effect of the explanary variables with better-defined and less-noisy patterns.